OpenStudy (anonymous):

Find the derivative of the function y = 4x^4 - 15x - (5/x^3) y' = 16x^3 - 15 and I'm confused on how to deal with (5/ x^3).

OpenStudy (anonymous):

5x^-3

OpenStudy (anonymous):

would that be clearer to deal with?

OpenStudy (anonymous):

apply division theorem used in diff

OpenStudy (anonymous):

5/(x^3) =5*x^(-3)

OpenStudy (anonymous):

or do as @mathsmind said

OpenStudy (anonymous):

recall nx^n-1

OpenStudy (anonymous):

y'=16x3-15-5(-3)/x4 =16x3-15+15/x4

OpenStudy (anonymous):

ah okay. so I I had something like 5 / 2x^3 it would be equal to 10x ^-3 then the derivative of that -30x ^-4?

OpenStudy (anonymous):

\[\frac{ 5 }{ 2 }x^{-3}\]

OpenStudy (anonymous):

try it again now...

OpenStudy (anonymous):

\[\frac{5}{2} (-3) x^{(-3-1)}\]

OpenStudy (anonymous):

-15/2(x)^-4

OpenStudy (anonymous):

can you visualize the pattern now?

OpenStudy (anonymous):

let 5/x^3 =5x^-3 =-3(5)x^-3-1 =-15x^-4

OpenStudy (anonymous):

I get that part I just didn't know how to deal with that fraction (5/x^3). Thanks for clarifying this.

OpenStudy (anonymous):

nx^(n-1), for n = -3